if p is jointly proportional to q and inversely proportional to s, state in words the relationship between q and the other variables
Answer by sachi(548)

if p is jointly proportional to q
i,e p=k1*q (k1=constant)
p is inversely proportional to s,
i.e. p=k2/s (k2=constant)
so p=k1*q*k2/s=k1*k2*q/s=kq/s (k1*k2=k=constant)